# A closer look at some cyclic semifields

### v1

@@ 6.1. $(f,\sigma,\delta)$-codes
for any two distinct generators $\sigma_{1}$ and $\sigma_{2}$ of the Galois group $\mathrm{Gal}(K/F)$ and any choice of $a_{i}\in K\smallsetminus F$, we expect that the choice of generator $\sigma_{i}$ will also affect the $\sigma_{i}$-constacyclic codes we obtain, since these correspond to the left ideals of two non-isomorphic algebras.
6.2. MRD-codes
Linear maximum rank distance codes (MRD-codes) can be obtained from skew polynomial rings $K[t;\sigma]$, employing the matrices representing the left multiplication in a division algebra $K[t;\sigma]/K[t;\sigma]f$, with entries in $K$.
Isotopy classes of finite semifields of dimension $n$ over $F=\mathbb{F}_{q}$ are in one-one correspondence with equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{n}(\mathbb{F}_{q})$ with minimum distance $n$. Over finite fields, Sandler showed that two algebras $K[t,\sigma]/K[t,\sigma](t^{m}-a)$ and $K[t,\sigma]/K[t,\sigma](t^{m}-b)$ are isotopic iff they are isomorphic.
Our results show that for $n\geq m$, the MRD codes obtained from the Petit algebras $K[t,\sigma]/K[t,\sigma](t^{m}-a)$ depend on the choice of the generator of the Galois group [She2019]: Different choices of generators of ${\rm Gal}(K/F)$ yield non-equivalent $F$-linear MRD-codes.
Since these codes are special cases of the codes constructed in [She2019] we conjecture that the equivalence classes of all of Sheekey’s codes also will depend on the choice of generator of the Galois group involved.
Moreover, we achieved an explicit parametrization of all semifields that are nonassociative cyclic algebras of a fixed prime degree over a fixed base field $F=\mathbb{F}_{q}$ which contains a primitive $n$th root of unity. That means we computed the number of non-isotopic semifields of order $q^{m^{2}}$ obtained via the construction $(K/F,\sigma,a)$ when ${\rm Fix}(\sigma)=F=\mathbb{F}_{q}$, and gave one representative for each isotopy class. Isotopy classes of finite semifields of dimension $m$ over $F=\mathbb{F}_{q}$ are in one-one correspondence with equivalence classes of $\mathbb{F}_{q}$-linear MRD-codes in $M_{m}(\mathbb{F}_{q})$ with minimum distance $m$. Thus we now have one representative for each corresponding MRD-code.

### v2

@@ 6.2. $(f,\sigma,\delta)$-codes
for any two distinct generators $\sigma_{1}$ and $\sigma_{2}$ of $\mathrm{Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$ and any choice of $a_{i}\in\mathbb{F}_{q^{n}}\smallsetminus\mathbb{F}_{q}$, the choice of generator $\sigma_{i}$ will also affect the $\sigma_{i}$-constacyclic codes we obtain, since these correspond to the left ideals of two non-isomorphic algebras.
First steps to classify skew $\sigma$-constacyclic codes by some natural $(m,\sigma)$-equivalence (respectively, some $(m,\sigma)$-isometry) that preserves their Hamming distance have recently been achieved in [Ou2025]. The main idea is that $\sigma$-constacyclic codes belonging to the same equivalence class have the same minimum distance and the same algebraic structures. The approach defines $(m,\sigma)$-equivalences/isometries using isomorphisms between Petit algebras that preserve the Hamming distance, but all codes are being built with the same fixed $\sigma$ that generates the Galois group $\mathrm{Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$.
The following definition refines this approach, and aligns with the definition of $(m,\sigma)$-equivalence/isometry given in [Ou2025]: let $a_{i}\in\mathbb{F}_{q^{n}}^{\times}$ and let $\sigma$, $\sigma_{j}$ be generators of $\mathrm{Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$. The skew $(\sigma_{i},a_{i})$-constacyclic codes correspond to ideals in $\mathbb{F}_{q^{n}}[t;\sigma_{i}]/\mathbb{F}_{q^{n}}[t;\sigma_{i}](t^{m}-a_{i})$ ($i=1,2$). Let $\mathcal{C}_{(\sigma_{i},a_{i})}$ be the class of skew $(\sigma_{i},a_{i})$-constacyclic codes related to $t^{m}-a_{i}\in\mathbb{F}_{q^{n}}[t;\sigma_{i}]$. Now $a_{1}$ and $a_{2}$ are called $(m,\sigma)$-isometric, if there exist an isomorphism between nonassociative Petit algebras
$\varphi:\mathbb{F}_{q^{n}}[t;\sigma_{1}]/\mathbb{F}_{q^{n}}[t;\sigma_{1}](t^{m}-a_{1})\longrightarrow\mathbb{F}_{q^{n}}[t;\sigma_{2}]/\mathbb{F}_{q^{n}}[t;\sigma_{2}](t^{m}-a_{2})$
that preserves the Hamming distance, and called $(m,\sigma)$-equivalent, if there exist $\alpha\in\mathbb{F}_{q}^{\times}$, such that
$\varphi_{\alpha}:\mathbb{F}_{q^{n}}[t;\sigma_{1}]/\mathbb{F}_{q^{n}}[t;\sigma_{1}](t^{m}-a_{1})\longrightarrow\mathbb{F}_{q^{n}}[t;\sigma_{2}]/\mathbb{F}_{q^{n}}[t;\sigma_{2}](t^{m}-a_{2}),\quad h(t)\longmapsto h(\alpha t)$
is an isomorphism between nonassociative Petit algebras (it might be less confusing to view these in this context simply as nonassociative rings). It is straighforward to see that if such an isomorphism $\varphi_{\alpha}$ exists then it will preserve the Hamming distance, as well as the algebraic structure by its nature, i.e. the ideals and their generators will be mapped $1-1$. Thus $\mathcal{C}_{(\sigma_{1},a_{1})}$ and $\mathcal{C}_{(\sigma_{2},a_{2})}$ will be equivalent classes.
But we know already that such isomorphisms do not exist when $n\geq m-1$ and $f=t^{m}-a$ is not right-invariant. We conclude:
Theorem 6.4.
Let $n\geq 3$, $n\geq m-1$, and let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$. Let $a_{i}\in K\smallsetminus F$, then $a_{1}$ and $a_{2}$ are not $(m,\sigma)$-equivalent and not $(m,\sigma)$-isometric.
Proof.
This follows immediately from $K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})\not\cong K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2}).$
It is still an open problem if this implies that the class of the skew $(\sigma_{1},a_{1})$-constacyclic codes and the one of the skew $(\sigma_{2},a_{2})$-constacyclic codes are not equivalent, either.